OpenStudy (anonymous):

find dy/dx if Y=u^1/3 and u=x^4-4x^3-4. I get 1/3(4x^7-28x^6+48x^5-16x^3+48x^2) This does not seem like it is correct to me.... What do you think?

OpenStudy (anonymous):

I have made a mess out of this one. I really need help....

OpenStudy (amistre64):

Chain rule: dy/dx = dy/du * du/dx dy/du= 1/3 * u^(-2/3) du/dx = 4x^3 - 12x^2 dy/dx = 4x^2(x - 3)/3(x^4 - 4x^3 -4)^(2/3) Looks about right to me :)